1. 6 + 5x = 2x- 2/3
Answer by mananth(16946)

1. 6 + 5x = 2x- 2/3
5x-2x= -2/3 - 1.6
3x=0.67-1.6
3x=-2.27
x=- -2.27/3
x=0.76